

********************************************************************
UARS WEEKLY STATUS REPORT - 26 FEB 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
data without apparent problems.
Reprocessing of the SUSIM data to data version 20 continues.
***********
PEM
***********
PEM is active under PMON control. Energy processor #6 for the
HEPS-1 unit has been found to not activate correctly. Data affected
are from the upper energy EE channel for Head-2 of HEPS-1 (that is the
+15 deg electrons above 300 keV). Tests are being conducted on this
energy processor to assure that it is operating correctly. Measurements
from this sensor will be activated as soon as possible. The PEM team
and FOT are convinced that the entanglement of energy processor #6
is related to OBC control and not a failure of that energy processor.
Early attempts at adjusting OBC command timing have not been successful.
The FOT is attempting to refine their diagnostics and confine the
problem.
While under PMON control, the PEM instrument computes energy deposition
from the HEPS unit.
***********
HALOE
***********
HALOE was turned on February 23, 1999. All engineering parameters are nominal.
***********
MLS
***********
The instrument is now off as scheduled. MLS is scheduled to begin south viewing
operations on Feb 27 and operate though Mar 4.
***********
HRDI
***********
HRDI continues to operate normally.
Recently, a significant recalibration of the HRDI interferometers was
performed. The HRDI Fabry-Perot etalons have been slowly drifting out of
parallelism since the beginning of the mission and this has led to a slow
decrease in the instrument sensitivity. On December 31, 1998, the HRDI
instrument performed a calibration to allow this drift to be corrected.
These calibration data were used to calculate new coefficients for the
instrument computer. On January 29, 1999 the new coefficients were entered
into the instrument processor and have been used by the instrument since.
As expected, the signal levels for all observed features rose to levels
comparable to those that were observed at the beginning of the mission. A
consequence of this change of coefficients is that the "zero wind" value
for the instrument is now different and currently unknown which introduces
a bias into the wind. We are in the process of evaluating the new zero,
but until that is completed the HRDI data products from January 29 onward
should not be used. We will notify the community when the data is again
valid.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
6 days ATT (7.3) RSS (Y2K Version)(6R-6C=0L)
0 days HAL (4.80) PROD (L3) (1874R-1721C-1F=152L)
2 days HAL (4.81) PROD (TEST L1-L3) (2R-2C=0L)
46 days PEM (3.32) PROD (OA-L2) (760R-130C=630L)
67 days SUS (7.0) PROD (647R-67C=580L)
The CDHF re-ingested level 0 for UARS day 2710 (julian 99/042)
due to better data quality. This ingest was completed on 02/18/99.
The CDHF is scheduling the upgrade of Multinet to V4.1b as part of the ongoing
Y2K implementation effort.
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres
DAAC data created in the month of December 1998 has been loaded into the
DAAC Transfer Tables for HALOE, HRDI, NMC, PEM, and UKMO.
***********
FOT
***********
This report covers 13 February 1999 (Orbit 40571, GMT Day 044 through 19
February 1999
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 048/2241 40645 Off
PEM 049/0017 40646 Cal test
The beta angle ranged between the angles of 34.1 degrees to 6.1 degrees for
this report period. The beta angle is now decreasing toward a minimum
angle of -2 degrees for the next yaw on 20 February 1999 (DOY 051).
The annular solar eclipse that occurred on 16 February 1999 over the Indian
Ocean, Australia and the South Pacific Ocean had little to no effect on
UARS power.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.3 degrees C. Additional battery performance data
for each battery on 12 February 1999 and 19 February 1999 are provided in
Appendix D.
The clock error ranged from 9.9 to -11.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 02
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 02
% HALOE 01
% PEM 03
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 05
Six (6) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/019 - 2287.496951
Transponder B: 99/020 - 2287.499045
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
Week Ending 12 February 1999
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/042 show the observatory data loss to be
22 hours, 46 minutes, 16.564 seconds (an increase of 55 minutes and 25.024
seconds since last report period). This is a 0.0351 percent data loss
which equals a 99.9649 percent data capture for the mission.
The increase of 55 minutes and 25.024 seconds occurred over two (2) days:
7 10.112 seconds on Day 036 due to data hit and insufficient overlap, and
7 55 minutes and 14.912 seconds on day 042 due to power outage
complications.
Week Ending 19 February 1999
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/049 show the observatory data loss to be
22 hours, 46 minutes, 25.972 seconds (an increase of 09.408 seconds since
last report period). This is a 0.0351 percent data loss which equals a
99.9649 percent data capture for the mission.
The increase of 09.408 seconds occurred over three (3) days:
7 0.512 seconds on Day 043 due to data hit,
7 8.384 seconds on Day 044 due to power outage complications, and
7 0.512 seconds on Day 047 due to data hit.
NCC switched their operational scheduling database to the "NCC-98" database
release on 12 February. Due to some minor problems with the migration of
"previously scheduled" TDRS contacts from the old database to the NCC-98
database, most POCCs were required to delete and re-add all of their
scheduled TDRS contacts spanning days 044-059. If the POCCs failed to
delete and re-add those events, the ability to send successful GCMRs during
the events in this time period would not exist. Due to the important need
for sending GCMRs, the UARS FOT deleted and re-added all of their
previously scheduled events during the day 044-059 time period.
The FOT received and processed the remaining Y2K "test" 8-week planning
aids that were requested by SOLSTICE operations (via the MPG). Two of the
seven planning aids were invalid and need to be regenerated by FDF. The
other five planning aids were copied to a special account on the UMPGAL
system, where they will be made available to all interested instrument
operations teams.
The UARS Test and Training Simulator (UTTS) remains non-operational. The
estimated time to repair the UTTS and return it to operation is unknown.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to support the PEM team with diagnosis of the Energy
Processor 6 status error. FOT attempted two tests, one a manual activation
of the instrument, and the other a redesign of the timing of PEM commands
for day-only. The first test established that EP6 can be commanded on by
the real-time activation sequence. The second test did not bring up EP6.
The FOT created another activation plan using real-time command sequences
to emulate the PMON activation sequence.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT - No change
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
Planning continues for the simulator testing.
UARS RE-ENGINEERING - TELEMETRY PROCESSING - No change
MEETINGS
FOT supported a meeting with Booz-Allen & Hamilton representatives (David
Taylor) on 18 February to discuss the daily operations timeline for UARS.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for 20 February 1999.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 19 February 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
40572 044/0142 99-066 OBC Clk Adj fr 0.00 to 20.6 msec/day
40595 045/1444 99-067 OBC Clk Adj fr 20.6 to 0.00 msec/day
40614 046/2052 99-068 PEM HEPS Command
40645 048/2241 99-069 HALOE Power OFF
40646 049/0017 99-070 PEM Calibration/testing
40662 050/0158 99-071 PEM Calibration/testing
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 19 February 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
NBRECBI Y-HI NBTR B Current 40612
NBRECBI Y-HI NBTR B Current 40627
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 19 February 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
40572 044/01:50 (12 Feb 99) from 0.00 to 20.6 msec/day
40595 045/14:48 (14 Feb 99) from 20.6 to 0.00 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 19 February 1999
12 Feb 1999 - GMT Day 043 Beta = 34.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 0.0/-11. -3.05/-4.31 +0.0/ +0.0 16.8 0.000 0.0
2 +56.0/ 0.0 +5.47/+4.51 +30.4/-14.4 26.2 1.030 21.1
3 +56.0/-11.2 +4.51/+3.55 +29.6/-14.4 26.1 1.027 20.8
19 Feb 1999 - GMT Day 050 Beta = 6.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 0.0/-22.4 -3.68/-4.94 +0.0/ +0.0 16.8 0.000 0.0
2 +67.2/ 0.0 +4.83/+3.87 +31.2/-14.4 25.6 1.016 20.8
3 +89.6/-5.6 +3.87/+2.60 +30.4/-14.4 25.4 1.017 20.4
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 19 February 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-007 FOT 40577 044/1019 N/A no lock on 32k OBC dump
99-008 FOT 40589 045/0419 20585 "No UARS RF, no TDEtrk"
99-009 FOT 40590 045/0650 14170 Generic Late Acquisition
99-010 FOT 40591 045/0838 N/A 6 masked ATC Dump errors
99-010 FOT 40636 048/0750 N/A GCMR to 32k on Q invalid
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-006 FOT see list Feb 99 14170 Generic Late Acq
SEVNTFW99/08